OpenStudy (ria23):

helllllllppppppp ; n ; 3x - y = -4 5x + 2y = 19 (Solve using elimination or substitution) I've tried this problem like, 500 times. I know that /SOMEHOW/ I'm supposed to come out with (1,7) but every time I do it... I end up with (11,-18) I have no idea what I'm doing wrong. >n<

OpenStudy (anonymous):

y=3x+4 5x + 2(3x+4) = 19 5x + 6x + 8 = 19 11x + 8 = 19 11x = 19 - 8 11x = 11 ... x = 1 y = 3(1) + 4 y = 7 so (x,y) ... (1,7)

OpenStudy (ria23):

Thank yhu so much for explaining. c:

OpenStudy (anonymous):

no worries